OpenStudy (anonymous):

A=2lw+2wh+2hl;solve for h

OpenStudy (hba):

Well you have to isolate h in this equation i guess. 1)Subtract both sides by 2lw.

OpenStudy (anonymous):

Because there is an 'h' in 2 different terms, you are going to have to get those terms alone on one side first by subtracting the other term from both sides. Once you have done that, you will have 2 terms on one side with a common factor of 'h'. Factor out that common factor. What do you get?

OpenStudy (anonymous):

so are you saying it would come to ...

OpenStudy (anonymous):

\[-2lw=2wh+2hl\]

OpenStudy (hba):

Where did the A go ?

OpenStudy (anonymous):

im sorry thats just bad math skills there i guess i removed it lol

OpenStudy (hba):

It would be, A-2lw=2wh+2hl Now take h common from right hand side. Can you do that ?

OpenStudy (anonymous):

A = 2lw + 2wh + 2hl -2lw - 2lw A - 2lw = 2wh + 2hl

OpenStudy (anonymous):

Similar example: 5xy - 5xz = x(5y - 5z)

OpenStudy (anonymous):

I factored out the common factor of x. You want to factor out the common factor of h on the right side and keep the left side.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

\[2wh+2hl=2(wh+hl)\]

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

2 was a common factor, but you were not interested in just any common factor. You wanted to get the h out of there. A - 2lw = h (2w+2l)

OpenStudy (anonymous):

now h is begin multiplied by (2w + 2l), so to get h alone, you have to do the opposite and divide both sides by (2w + 2l). Leave your answer as a fraction.

OpenStudy (anonymous):

a-2lw all over (2w+2l)=h

OpenStudy (hba):

Yeah right :) \[\frac{ A-2lw }{ 2w+2l }=h\]

OpenStudy (anonymous):

thank you @hba